OpenStudy (anonymous):

Geometry Question I Need To Find The Value Of X, Then Round to The Nearest Tenth. I Would Appreciate it if someone can show me how so solve this step by step. I Have a bunch of problems like this so i just need to learn how to do 1 and ill be on my way thanks in advance.

OpenStudy (anonymous):

I'll help :)

OpenStudy (anonymous):

Only #17 please

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

okay so first do you have a calculator with you ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay do you know about sin, cos, tan?

OpenStudy (anonymous):

Kind of ill bare with you

OpenStudy (anonymous):

okay good I will briefly explain how your going to use them in these problems

OpenStudy (anonymous):

so, first come the formulas which are....

OpenStudy (anonymous):

Sine = Opposite / Hypotenuse Cosine = Adjacent / Hypotenuse Tangent = Opposite / Adjacent

OpenStudy (ribhu):

100 sin inverse 20 deg = 34.20

OpenStudy (anonymous):

okay awesome!! so you know your formulas now lets put them to work

OpenStudy (anonymous):

so, let's look at #17

OpenStudy (anonymous):

sin20= x/100 from calculator/ table 0.34202 = x/100 x = 34.20

OpenStudy (anonymous):

lol exactly

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

your so welcome :)

OpenStudy (ribhu):